OpenStudy (anonymous):

what is the point slope of the line with slope -10 that goes thru (1, 4) ? A. y-4 = -10(x-1) B. y+4 = 10(x+1) C. y+4 = -10(x+1) D. y-4 = 10(x-1)

OpenStudy (anonymous):

Substitute the values and solve x=1 y=4, the one that proves itself is the answer

OpenStudy (anonymous):

or you could solve for the intercept points and graph--but that is a lot longer

OpenStudy (anonymous):

Point slope form is y-y1=m(x-x1) (x1,y1) is your point, m is the slope

OpenStudy (anonymous):

so...its C. y+4 = -10(x+1)!

OpenStudy (anonymous):

No. It's A, the sign changes on the points